OpenStudy (anonymous):

"Rationalize the denominator of each expression. Assume that all variables are positive." Problem: Square root of y/ square root of 5

OpenStudy (anonymous):

To rationalize denominators which just include a single sqrt you would multiply the numerator and denominator by the denominator's sqrt. This will remove the radical from the denominator. For example: \[\frac{ \sqrt{9x} }{ \sqrt{15} }*\frac{ \sqrt{15} }{ \sqrt{15} }=\frac{ 3\sqrt{15x} }{ 15}=\frac{ \sqrt{15x} }{ 5 }\]

OpenStudy (anonymous):

what if it is a cubed root? would it be the same? \[\sqrt[3]{7xy^2}\div \sqrt[3]{4x^2}\]

OpenStudy (anonymous):

I would probably do the second one differently. I would use rules of rational exponents. See my solution steps in your other post.